Hello, this is all one question. Thank you very much in advance! They want me to fill in the blanks! Thank you!

BALANCE SHEET ANALYSIS

Complete the balance sheet and sales information using the following financial data:

Total assets turnover: 1.6x
Days sales outstanding: 30 daysa
Inventory turnover ratio: 4x
Fixed assets turnover: 3.5x
Current ratio: 2x
Gross profit margin on sales: (Sales - Cost of goods sold)/Sales = 25%
aCalculation is based on a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.

Balance Sheet
Cash $   Current liabilities $  
Accounts receivable    Long-term debt 60,000
Inventories    Common stock   
Fixed assets    Retained earnings 100,000
Total assets $400,000 Total liabilities and equity $  
Sales $   Cost of goods sold $  
Answer #1

Total assets turnover = Sales / Total Assets

1.6 = Sales / $400,000

Sales = 1.6 x $400,00

Net Sales = $640,000

Gross Profit Ratio = 25%

(Sales - Cost of goods sold) / Sales = 0.25

($640,000 - Cost of goods sold) / $640,000 = 0.25

$640,000 - Cost of goods sold = $160,000

Cost of goods sold = $640,000 - $160,000

Cost of goods sold = $480,000

Days sales outstanding: 30 days

Accounts Receivable = (30 / 365) * Sales

Accounts Receivable = (30 / 365) * $640,000

Accounts Receivable = $52,603

Inventory turnover ratio: 4x

Inventory turnover ratio = Cost of Goods sold / Inventory

4 = $480,000 / Inventories

Inventories = $480,000 / 4

Inventories = $120,000

Fixed assets turnover: 3.5x

Fixed assets turnover ratio = Sales / Fixed Assets

3.5 = $640,000 / Fixed Assets

Fixed Assets = $640,000 / 3.5

Fixed Assets = $182,857

Total Assets = $400,000

$400,000 = Cash + Accounts receivable + Inventories + Fixed assets

$400,000 = Cash + 52,603 + 120,000 + 182,857

Cash = $400,000 - (52,603 + 120,000 + 182,857)

Cash = $400,000 - $355,460

Cash = $44,540

Current ratio: 2x

Current ratio = Current Assets / Current Liabilities

Current Assets = Cash + Accounts receivable + Inventories

2 = (44,540 + 52,603 + 120,000 ) / Current Liabilities

2 = 285,714 / Current Liabilities

Current Liabilities = 217,143 / 2

Current Liabilities = $108,572

We know, Total Assets = Total liabilities and equity

That means, Total liabilities and equity = $400,000

Total liabilities and equity = Current liabilities + Long-term debt + Common stock + Retained earnings

$400,000 =  108,572 + 60,000 + Common stock + 100,000

$400,000 = 268,572 + Common stock

Common stock = 131,428

Balance Sheet $ $ Current Liablities Cash 44,540 108,572 Accounts Receivable Long-term debt 52,603 60,000 Common Stock 131,42

If any doubts ask me in the comments.

If you like my answer give me a thumbs up! Thank you!